find the length of the diagonal of a rectangle whose sides are radical 5 inches and 2 inches respectively ? please show all work
Answer by checkley77(12844)

A^2+B^2=C^2
(SQRT5)^2+2^2=C^2
5+4=C^2
C^2=9
C=SQRT9
C=3 INCHES IS THE DIAGONAL.